ADNAN MUHAMMED

Grade 12,

A block of mass 2.00 kg moving at speed of 10.0 m/s accelerates at 3.00 m/ss for 5.00 s.compute its final kinetic energy.

A block of mass 2.00 kg moving at speed of 10.0 m/s accelerates at 3.00 m/ss for 5.00 s.compute its final kinetic energy.

Grade:12th pass

1 Answers

Vikas TU
7 years ago
After accelerating the final velocity becomes:
v = 10 + 3*5 = >10 + 15 = 25 m/s.
Hence its  final Kinetic Energy becomes = > 0.5mv^2 = > 0.5 * 2 * 25*25 => 625 Joule.